# Physics-Informed Dynamical Modeling of Extrusion-Based 3D Printing Processes

### v1

@@ 4.3 Model evaluation under different surface velocity conditions
In this subsection, we first examine the model’s interpolative performance with respect to surface velocity. The model is trained using datasets representing the lowest and highest surface velocities in the simulation suite, 50 mm/s and 70 mm/s, capturing the full range of nozzle–substrate relative motion. Training on these bounding cases exposes the model to both minimal and maximal shear and deformation effects imposed by the substrate motion. The trained model is subsequently tested on the intermediate surface velocity of 60 mm/s, which is excluded from the training data. This configuration allows us to evaluate how effectively the model predicts flow behavior for surface velocities that fall within the interior of the parameter space.
After interpolative scenario is evaluated, we investigate the model’s extrapolative performance with respect to surface velocity. In the first scenario, the model is trained on the intermediate and highest surface velocities, 60 mm/s and 70 mm/s, which capture moderate to high relative motion between the nozzle and the build plate. These datasets provide the model with the dominant deformation and shear patterns that arise near the upper end of the operating range. The model is then tested on the lowest surface velocity of 50 mm/s, which lies outside the training region. This allows us to assess how well the model generalizes when confronted with deposition conditions characterized by reduced substrate motion and lower shear rates.
Figure 11: RMSE distribution of the dynamical model across all the datasets and sub-systems for different interpolative and extrapolative $U_{s}$ scenarios. (a) The interpolative scenario is trained with cases 1, 3, 4, 6, 7, and 9 as the training cases and tested on cases 2, 5, and 8. (b) The low-end extrapolative scenario is trained with cases 2, 3, 5, 6, 8, and 9 as the training cases and tested on cases 1, 4, and 7. (b) The high-end extrapolative scenario is trained with cases 1, 2, 4, 5, 7, and 8 as the training cases and tested on cases 3, 6, and 9. Dark blue denotes the training cases, whereas light blue denotes the testing cases.
@@ 4.4 Model evaluation under randomized train–test partitioning
Figure 12: RMSE distribution of the dynamical model across all the datasets and sub-systems for random (a) 67-33, (b) 54-46, (c) 46-54, and (d) 33-67 training-testing splits.
@@
Across Figures 12(b) to (d), in general, case 3 exhibits the highest error levels for sub-system 2, expect for Figure 12(d) where the training data is scarce. This behavior is physically expected, as case 3 corresponds to the lowest inlet mass-flow-rate condition combined with a relatively high plate velocity. Under these circumstances, the liquid film becomes more strongly influenced by shear-driven motion, leading to a thinner and more highly stretched film. Such flow regimes are more sensitive to small variations in operating conditions and therefore produce stronger nonlinear coupling between inertia, viscous effects, and wall shear.

### v2

@@ 4.3 Model evaluation under different surface velocity conditions
In this subsection, we first examine the model’s interpolative performance with respect to surface velocity. The model is trained using datasets representing the lowest and highest surface velocities in the simulation suite, at 50 mm/s and 70 mm/s, capturing the full range of nozzle–substrate relative motion. Training on these bounding cases exposes the model to both minimal and maximal shear and deformation effects imposed by the substrate motion. The trained model is subsequently tested on the intermediate surface velocity of 60 mm/s, which is excluded from the training data. This configuration allows us to evaluate how effectively the model predicts flow behavior for surface velocities that fall within the interior of the parameter space.
After interpolative scenario is evaluated, we investigate the model’s extrapolative performance with respect to surface velocity. In the first scenario, the model is trained on the intermediate and highest surface velocities, at 60 mm/s and 70 mm/s, which capture moderate to high relative motion between the nozzle and the build plate. These datasets provide the model with the dominant deformation and shear patterns that arise near the upper end of the operating range. The model is then tested on the lowest surface velocity of 50 mm/s, which lies outside the training region. This allows us to assess how well the model generalizes when confronted with deposition conditions characterized by reduced substrate motion and lower shear rates.
@@
Figure 11: RMSE distribution of the dynamical model across all the datasets and sub-systems for different interpolative and extrapolative $U_{s}$ scenarios. (a) The interpolative scenario is trained with cases 1, 3, 4, 6, 7, and 9 as the training cases and tested on cases 2, 5, and 8. (b) The low-end extrapolative scenario is trained with cases 2, 3, 5, 6, 8, and 9 as the training cases and tested on cases 1, 4, and 7. (c) The high-end extrapolative scenario is trained with cases 1, 2, 4, 5, 7, and 8 as the training cases and tested on cases 3, 6, and 9. Dark blue denotes the training cases, whereas light blue denotes the testing cases.
@@ 4.4 Model evaluation under randomized train–test partitioning
Figure 12: RMSE distribution of the dynamical model across all the datasets and sub-systems for random (a) 67-33, (b) 54-46, (c) 46-54, and (d) 33-67 training-testing splits. Dark blue denotes the training cases, whereas light blue denotes the testing cases.
@@
Across Figures 12(b) to (d), in general, case 3 exhibits the highest error levels for sub-system 2, except for Figure 12(d), where the training data are scarce. This behavior is physically expected, as case 3 corresponds to the lowest inlet mass-flow-rate condition combined with a relatively high plate velocity. Under these circumstances, the liquid film becomes more strongly influenced by shear-driven motion, leading to a thinner and more highly stretched film. Such flow regimes are more sensitive to small variations in operating conditions and therefore produce stronger nonlinear coupling between inertia, viscous effects, and wall shear.